# UniFoil: A Universal Dataset of Airfoils in Transitional and Turbulent Regimes for Subsonic and Transonic Flows

### v2

@@ Abstract
We present UniFoil, the largest publicly available universal airfoil dataset based on Reynolds-averaged-Navier-Stokes (RANS) simulations. It contains over 500,000 samples spanning a wide range of Reynolds numbers, Mach numbers, capturing both transitional and fully turbulent flows across incompressible to compressible regimes. UniFoil is designed to support machine learning (ML) research in fluid dynamics, particularly for modeling complex aerodynamic phenomena. Most existing datasets are limited to incompressible, fully turbulent flows with smooth field characteristics, thus overlooking the critical physics of laminar–turbulent transition and shock-wave interactions-features that exhibit strong nonlinearity and sharp gradients. UniFoil fills this gap by offering a broad spectrum of realistic flow conditions. Turbulent simulations utilize the Spalart–Allmaras (SA) model, while transitional flows are modeled using an $e^{N}$-based transition prediction method coupled with the SA model. The dataset includes a comprehensive geometry set comprising over 4,800 natural laminar flow (NLF) airfoils and 30,000 fully turbulent (FT) airfoils, effectively covering the diversity of airfoil designs relevant to aerospace, wind energy, and marine applications. This dataset is also highly valuable for scientific machine learning (SciML), enabling the development of data-driven models that more accurately capture the transport processes associated with laminar–turbulent transition. UniFoil is freely available under a permissive CC-BY-SA license.
@@ 1 Introduction
The remainder of this paper is organized as follows. In Section 1, we review existing datasets and highlight the unique features of our proposed dataset, UniFoil. Section 2 details the tools and techniques used for data generation, while Section 3 describes the dataset structure and access details. Finally, conclusions are presented in Section 4.
@@ 1.1 Related Work
There are several airfoil datasets represent significant advancements in providing large-scale, ML-ready CFD data tailored for aerodynamic analysis. The datasets information is summarized in Table 1. The first dataset by Schillaci et al. (2021), comprises 425 simulations of turbulent flows over airfoils without modeling laminar–turbulent transition. The simulations are incompressible, and volume-resolved field data are provided for each case. The second dataset, AIRFRANS Bonnet et al. (2022), includes 1,000 simulations of turbulent airfoil flows at a Mach number (Ma) of 0.3 and Reynolds number (Re) ranging from 2 to 6 million. The dataset does not model transition and is limited to incompressible flows. It provides full field data across a wide range of angles of attack (AoAs). The third dataset by Ramos et al. (2023a), expands to 1,830 airfoils and incorporates both turbulence and transition modeling. Although the flows remain incompressible with a fixed Mach number of 0.1, the dataset spans multiple Re (3, 6, and 9 million) and AoAs from $-4^{\circ}$ to $20^{\circ}$ and includes over 250,000 simulations with corresponding flow field data. Other similar datasets have been summarized in Table 1 and compared with our UniFoil dataset.
@@
Dataset No. of Airfoils Turb. Trans. Model Compr. Mach Re (×$\times 10^{6}$) AoA () Field Data No. of Sim.
@@ 1.2 Focus of this work
The existing dataset focuses on incompressible fully turbulent flows which missed critical strongly nonlinear physics including shock waves and laminar turbulent transition. While our new dataset is more general with these essential physics incorporated. More details can be found in Figure 1.
A shock wave is a narrow region across which flow properties such as pressure, temperature, and velocity change abruptly due to compressibility effects at high speeds (see the second figure in the first row of Figure 1). Shock waves introduce sharp spatial discontinuities in fluid flow that pose fundamental challenges for ML models, particularly those relying on smoothness assumptions or local interpolation. In high-speed regimes, shocks dominate the flow field and drive nonlinear phenomena such as boundary layer separation and wave interactions. For ML researchers, modeling and predicting shocks tests the limits of both inductive bias and network expressivity, especially in generalization across flow regimes. By including shock-resolving simulations in our dataset, we offer a benchmark that pushes beyond existing datasets constrained to subsonic, smooth flows—highlighting scenarios where naive ML architectures fail and physics-aware models are essential.
@@ 1.3 Main contribution
To summarize and contrast the key new additions the current dataset brings to the literature, we refer the readers to Table 1 and a summary of the input fields can be found in Table 2.
@@ 2.1 Geometry
Airfoil geometries can be efficiently represented in a reduced state space using modal decomposition techniques. In particular, a low-dimensional basis can be extracted from a dataset of classic airfoils via proper orthogonal decomposition (POD), enabling compact representation and efficient sampling of shape variations. Notably, airfoils designed for fully turbulent operation differ significantly in geometry—and consequently in modal characteristics—from those optimized to exploit laminar–turbulent transition mechanisms. Thus, we create two airfoil shape datasets for the fully turbulent (FT) airfoils and natural laminar flow (NLF) airfoils. In the current dataset, we sampled 30,000 FT and 4,800 NLF airfoils. An example of this sampling is shown in Figure 2 for the FT airfoils. We refer the readers to Li et al. (2019) for further information on sampling large datasets of geometric variations in a particular class of airfoils.
@@ 2.2 Mesh generation
Mesh generation for the airfoil simulations was carried out using pyHyp by Secco et al. (2021) [^2] https://github.com/mdolab/pyhyp. pyHyp generates high-quality O-type multi-block meshes around airfoils using a hyperbolic marching method, which extrudes mesh lines from an initial surface mesh by solving a system of hyperbolic partial differential equations (PDEs) that control cell spacing, orthogonality, and smoothness. One mesh generated by pyHyp is shown in Figure 3/ This approach enables accurate boundary layer resolution and consistent mesh quality, making it well suited for RANS solvers such as ADflow. The method is fully automated and robust across large geometric datasets, supporting high-throughput CFD simulations.
@@ 2.3 Solver and numerical considerations
All simulations were steady state and performed on a far-field mesh with the airfoil body-fitted at the center of the mesh, as shown in Figure 3. The wall was given a no-slip boundary condition and the outermost edge of the computational domain was given a far-field boundary condition. Further details on the numerical methods employed in data generation are elaborated in the supplementary data.
An overview of the flow field figures and flow regimes in the dataset are summarized in Figure 1. In the figure, $C_{p}$ is the coefficient of pressure and $\mu_{t}/\mu$ is the ratio of the turbulent eddy viscosity to the fluid’s molecular viscosity.
@@ 3 Dataset description
We sample a wide range of Mach numbers, Reynolds numbers, and AoAs. We have the largest ranges of Mach numbers and Reynolds numbers among all datasets Table 1. The range of AoA is narrower than in many other datasets, as large AoAs are typically avoided in practice to maintain aerodynamic efficiency. In addition, with a larger AoA, RANS model is no longer able to capture the flow field accurately due to flow separation. For each airfoil geometry, we conduct around $13$ simulations with different Mach, Re, and AoA. These three input variables were sampled using a Lattice HyperCube Sampling approach.
The dataset is composed of three parts: The majority of the simulations are done using the fully turbulent flow assumption with SA model together with fully turbulent airfoil geometry class (more details see Section 2). This dataset represents the most commonly used operation condition. The second dataset also uses the fully turbulent assumption but uses the NLF geometry class. Finally, the last dataset has a one-to-one correspondance with the second dataset with identical flight conditions and geometries. The only difference is that the last dataset employs an $e^{N}$-based transition model to capture the laminar–turbulent transition, whereas the second dataset assumes fully turbulent flow from the leading edge of the airfoil.
@@
Turbulence Model Airfoil Class No. of Airfoils Mach Re (×$\times 10^{6}$) AoA () Net Simulations
SA model FT 30,000 $[0.1,0.85]$ $[1,10]$ $[-2,6]$ 400,000
NLF 4,800 $[0.1,0.85]$ $[1,10]$ $[-2,6]$ 50,000
$e^{N}$ transition model (+ trip term) NLF 4,800 $[0.1,0.85]$ $[1,10]$ $[-2,6]$ 50,000
@@ 3.1 Dataset access and contents summary
The dataset is publicly available under the CC-BY-SA license at the Harvard Dataverse Kanchi et al. (2025). The dataset contents are summarized in Table 3. All data processing and visualization scripts will be released in a public repository referenced in the supplementary material. The CGNS files listed in Table 3 have the coefficient of pressure, velocity vector data and the mach number. Each CGNS file retains the grid and the values of the computed flow variables in that grid. For the NLF-transition simulations, the CGNS files have additional information, such as turbulence intermittency, temperature, skin-friction coefficients and density. The lift and drag coefficients ($C_{L}$ and $C_{D}$) are placed in the tabular file as described in Table 3.
@@
airfoil_geometry/ Airfoil coordinates and generation code (in supplementary repo).
@@ 3.3 Limitations of the Dataset
4 Conclusion
5 Acknowledgement
Appendix A Checklist
This section completes the checklist for the Neurips 2025 Datasets track submission.
  • Claims:
    Response - Yes. The data presented in the manuscript in preceding sections suffices for the support of the paper’s contributions and its scope.
  • Limitations:
    Response - Yes. We have an explicit section (Section 3.3) in the paper that details the limitations of the dataset and re-inforces the scope of the manuscript.
  • Theory, Assumptions and Proofs:
    Response - NA. The paper does not portray any theoretical results.
  • Experimental Result Reproducibility:
    Response - Yes. The paper has a supplementary data section that provides details on how the data can be generated. The tools used in the dataset generation are all open-source.
  • Open Access to Data and Code:
    Response - Yes. All data is open-access.
  • Experimental Setting/ Details:
    Response - Yes. Details on the dataset structure and the input parameters to all simulations have been clearly provided in this manuscript in a file in the dataset.
  • Experiment Statistical Significance:
    Response - NA.
  • Experiments Compute Resource:
    Response - Yes. A section (Section 3.2) has been made to address this query.
  • Code Of Ethics:
    Response - Yes.
  • Broader Impacts:
    Response - NA.
  • Safegaurds:
    Response - NA.
  • Licenses:
    Response - Yes. All the work is published and available under the CC-BY-SA license.
  • Assets:
    Response - NA.
  • Crowdsourcing and Research with Human Subjects:
    Response - NA.
  • IRB Approvals
    Response - NA
  • Declaration of LLM usage
    Response - NA
Appendix B Geometry sampling
$\mathbf{Y}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\mathrm{T}},$ (2)
$\widetilde{\mathbf{U}}\in\mathbb{R}^{n\times r},$ (3)
$\mathbf{y}=\widetilde{\mathbf{U}}\tilde{\mathbf{y}},$ (4)
The choice of the modal-coefficients is key in generation of the airfoil shape database. Thus, instead of representing each airfoil with the ($x,y$) coordinate values of its profile, we use the reduced few modal coefficients that were used to generate that airfoil in the input data to the neural network. For more details on choice of modal coefficients and how the sampling is done exactly for large number of airfoils, we direct the readers to Li et al. [2019, 2020]. In this manner, the $4,800$ NLF airfoils and $30,000$ FT airfoils were sampled. The code to sampling these airfoils and list of modal coefficients used for the generation of the dataset are detailed in the GitHub repository publicly available in the CC-BY-SA license [^3] GitHub Repository - https://github.com/rohitroxkp7/UniFoil.
Appendix C Numerical method
C.1 Governing equations
The turbulence source term is defined as follows
$\bm{s}(\bm{q})=\begin{bmatrix}0\\ \bm{0}\\ 0\\ C_{b1}(1-f_{t2})\bar{S}\tilde{\nu}+\frac{C_{b2}}{\sigma}\absolutevalue{\nabla\tilde{\nu}}^{2}-C_{w1}f_{w}\left(\frac{\tilde{\nu}}{d}\right)^{2}\end{bmatrix},$ (7)
$g\left(C_{b1}(1-f_{t2})\bar{S}\tilde{\nu}+\frac{C_{b2}}{\sigma}\absolutevalue{\nabla\tilde{\nu}}^{2}-C_{w1}f_{w}\left(\frac{\tilde{\nu}}{d}\right)^{2}\right).$ (9)
C.2 Grid generation
We use pyHyp Secco et al. [2021] to generate the volume grid, which uses a hyperbolic volume grid marching scheme to extrude structured surface grids into volume grids. Thus, the grid generation was performed such that the domain is an O-type grid with the airfoil body–fitted to the center of the domain with one cell thickness in the $z$-direction as shown in Figure 3.
Table 4: Fine grid generation parameters
Description Value
Number of points along airfoil surface till TE 293
Number of TE points 11
Number of points in marching direction 85
First layer thickness $10^{-6}$ c
Distance from the airfoil surface to march the hyperbolic grid $100c$
To maintain the same shape for the input data format, we ensured to maintain the same input parameters for the grid generation. The input parameters are summarized in Table 4. In the table, “TE” stands for trailing edge and “$c$” stands for chord, which was set to a value of 1m in all simulations.
Appendix D Output file contents
As introduced in Section 3, the dataset comprises three categories of simulations:
  1. 1.
    Fully turbulent simulations for FT airfoils,
  2. 2.
    Fully turbulent simulations for NLF airfoils,
  3. 3.
    Transition simulations for NLF airfoils.
For each of the output files, we describe the contents in the following subsections. All simulation output files follow a consistent naming convention and are provided in CGNS format. The structure and content of these files are summarized in Table 5. In the following subsections, we discuss the output folders and files content into two sub–categories based on the geometry, the first being the FT–airfoil dataset and second being the NLF airfoil dataset.
Governing equation Airfoil Folder Suffix Freestream File
D.1 FT–airfoil dataset
Turb_Cutout_<1--6>.zip.
Each simulation file follows the naming pattern:
Airfoil_Case_Data_turb.tab .
A visualization of the field data present in the fully–turbulent dataset is shown in Figure 4. In the figure, we show the Mach number, velocity magnitude and pressure coefficient values. It also showcases shocks on both the suction and pressure sides, with flow separation aft the shock foot. The velocity magnitude $|\mathbf{u}|$ is computed using the formula
$|\mathbf{u}|=\sqrt{u^{2}+v^{2}},$ (10)
D.2 NLF–airfoil dataset
NLF_Airfoils_Fully_Turbulent.zip ,
with each simulation file inside this folder that follows the name:
Airfoil_Case_Data_Trans_lam.tab ,
for both the fully-turbulent and transition datasets of the NLF–airfoils. Each CGNS file, once again, contains the Mach number, pressure coefficient, $C_{p}$, and the velocity vector field, as shown in Figure 4 for the FT-airfoils.
airfoil_<airfoil_num>_G2_A_L0_case_<case_num> .
Each of these subfolders contains the following files:
Finally, the archive Transi_sup_data_cutout_<1--4>.zip mirrors this folder structure and includes eight supplementary files per case. These provide extended data on the transition characteristics, enabling deeper analysis of the laminar-to-turbulent transition region. A full description of these supplementary files and their contents is available in the accompanying README file on the GitHub repository [^4] GitHub Repository - https://github.com/rohitroxkp7/UniFoil.
A visualization of the velocity magnitude field and the field of ratio of turbulent eddy viscosity to the molecular viscosity is shown in Figure 5. The transition onset points are marked on the airfoil surface using the red–colored arrow on the suction side of the airfoil. The transition to turbulence was only observed for the suction side in this case. The location for transition point was obtained from the “transiLoc.dat” file.
In the $e^{N}$ transition model, the reduced TS wave frequency is given by
$f=\frac{\omega}{Re_{\delta}},$ (11)
$N=\int_{x_{0}}^{x}\alpha_{i}(x,f)\,\text{d}x,$ (12)
${N}_{\text{TS}}=\text{max}_{f}(N(f)).$ (13)
Out of the eight files, the key files are “nfactor_ts.dat” and “transiLoc.dat”. The former has the TS wave factor $N_{\text{TS}}$ and the latter has the transition point location, if detected, on the suction and pressure surfaces. Details on accessing and visualizing these files can be found in the GitHub repository [^5] GitHub Repository - https://github.com/rohitroxkp7/UniFoil. Details on the transition characteristics and the factors can be found in the paper by Shi et al. [2020].
Figures of the skin-friction coefficient $C_{f}$ and the $N_{\text{TS}}$ factor plotted against $x/c$ from the nfactor_ts.dat file is shown in Figure 6, for the airfoil shown in Figure 5. The $N_{\text{TS}}$ factor for TS waves is computed as shown in Equation 13. When transition occurs, the skin–friction coefficient starts to abruptly increase. The transition region highlighted in Figure 6 can be seen for the suction–side of the airfoil where the skin friction coefficient suddenly rises to a value of roughly 0.0043 before it starts to dip. For the $N_{\text{TS}}$ factor, values begin to increase around $x/c$ of 0.2, in–line with the observation from the skin–friction coefficient plot.
Appendix E Data preprocessing
To ensure a smooth handling of input data, maintaining the same input shape to the neural network and to preserve the grid resolution without loss of data that typically occurs when interpolating onto a uniform grid, the data from the O-grid was reshaped into a neat 2D array. The process of reshaping is highlighted in Figure 7. We start at the TE and move clockwise along the airfoil surface to loop till we reach back to the TE. This forms the first bottom–most layer of the 2D array. We then move to the next layer and repeat till all the layers are completed.
Since the same tool (pyHyp) and grid generation inputs from Table 4 were used, we maintained the same number of “O-grid rings” and ordering of points to reshape all input data in a similar manner. Figure 7 shows this concept in the top row and shows the implementation of the reshaping algorithm for the velocity field in the bottom row.
Appendix F Preliminary results for Machine Learning
A neural network was trained using the UniFoil dataset. In this section, we discuss results from the training of this neural network. First, we discuss the details on network architecture. This includes details such as neuron activations, number of layers and the neural network setup. Then, we discuss the results from the neural network by comparing the reconstructed values against the ground–truth data.
F.1 Neural network architecture
The neural network trained on the UniFoil dataset is an encoder-decoder model with a latent space deep neural network (DNN). This model is trained on the pressure-coefficient data extracted from each airfoil simulation. The model is split into two main components with different layer designs used to down-sample and up-sample the input space.
An overview of the architecture of the neural network is shown in Figure 8. A summary of the training setup is shown in Table 6. In the table the “Param #” field represents the quantity of variables that can be modified during the training procedure. Densely connected layers such as the ones prevalent in the DNN have a large number of these parameters as all 6134 nodes are connected to every single node below.
The model was trained by first preparing the encoder-decoder model and training it on the reshaped pressure coefficent data. Once this model is trained, the encoder is used to convert all of the pressure coefficent data into latent space values which are saved seperately. These latent space values are then combined with matching airfoil feature data in order to train the deep neural network. When the deep neural network is done training, the model can be used to predict an airfoils pressure coefficent field by first converting the feature data into the latent space and then passing this latent space data into the decoder. This results in a pressure coefficient field in the square grid format which must finally be reshaped into the O-Grid using the geometry data from the airfoil.
Figure 8: An overview of the entire training pipeline. The modal coefficients $a_{i}$ are obtained from Equation 4.

### v4

@@ Abstract
We present UniFoil, the largest publicly available universal airfoil dataset based on Reynolds-averaged-Navier–Stokes (RANS) simulations. It contains over 500,000 samples spanning a wide range of Reynolds numbers, Mach numbers, capturing both transitional and fully turbulent flows across incompressible to compressible regimes. UniFoil is designed to support machine learning (ML) research in fluid dynamics, particularly for modeling complex aerodynamic phenomena. Most existing datasets are limited to incompressible, fully turbulent flows with smooth field characteristics, thus overlooking the critical physics of laminar–turbulent transition and shock-wave interactions-features that exhibit strong nonlinearity and sharp gradients. UniFoil fills this gap by offering a broad spectrum of realistic flow conditions. Turbulent simulations utilize the Spalart–Allmaras (SA) model, while transitional flows are modeled using an $e^{N}$-based transition prediction method coupled with the SA model. The dataset includes a comprehensive geometry set comprising over 4,800 natural laminar flow (NLF) airfoils and 30,000 fully turbulent (FT) airfoils, effectively covering the diversity of airfoil designs relevant to aerospace, wind energy, and marine applications. This dataset is also highly valuable for scientific machine learning (SciML), enabling the development of data-driven models that more accurately capture the transport processes associated with laminar–turbulent transition. UniFoil is freely available under a permissive CC-BY-SA license.
@@ 1 Introduction
The remainder of this paper is organized as follows. In Section˜1, we review existing datasets and highlight the unique features of our proposed dataset, UniFoil. Section˜2 details the tools and techniques used for data generation, while Section˜3 describes the dataset structure and access details. Finally, conclusions are presented in Section˜5. The dataset is available through the Harvard dataverse [^1] Harvard dataverse link - https://doi.org/10.7910/DVN/VQGWC4. Tools to help view the dataset with preliminary ML results are available in the GitHub repository [^2] GitHub Repository — https://github.com/rohitroxkp7/UniFoil.
@@ 1.1 Related Work
There are several airfoil datasets represent significant advancements in providing large-scale, ML-ready CFD data tailored for aerodynamic analysis. The datasets information is summarized in Table˜1. The first dataset by Schillaci et al. (2021), comprises 425 simulations of turbulent flows over airfoils without modeling laminar–turbulent transition. The simulations are incompressible, and volume-resolved field data are provided for each case. The second dataset, AIRFRANS Bonnet et al. (2022), includes 1,000 simulations of turbulent airfoil flows at a Mach number (Ma) of 0.3 and Reynolds number (Re) ranging from 2 to 6 million. The dataset does not model transition and is limited to incompressible flows. It provides full field data across a wide range of angles of attack (AoAs). The third dataset by Ramos et al. (2023a), expands to 1,830 airfoils and incorporates both turbulence and transition modeling. Although the flows remain incompressible with a fixed Mach number of 0.1, the dataset spans multiple Re (3, 6, and 9 million) and AoAs from $4^{\circ}$ to $20^{\circ}$ and includes over 250,000 simulations with corresponding flow field data. Other similar datasets have been summarized in Table˜1 and compared with our UniFoil dataset.
@@
Dataset No. of Airfoils Turb. Trans. Model Compr. Mach Re ($\times 10^{6}$) AoA () Field Data No. of Sim.
@@ 1.2 Focus of this work
The existing dataset focuses on incompressible fully turbulent flows which missed critical strongly nonlinear physics including shock waves and laminar turbulent transition. While our new dataset is more general with these essential physics incorporated. More details can be found in Figure˜1.
A shock wave is a narrow region across which flow properties such as pressure, temperature, and velocity change abruptly due to compressibility effects at high speeds (see the second figure in the first row of Figure˜1). Shock waves introduce sharp spatial discontinuities in fluid flow that pose fundamental challenges for ML models, particularly those relying on smoothness assumptions or local interpolation. In high-speed regimes, shocks dominate the flow field and drive nonlinear phenomena such as boundary layer separation and wave interactions. For ML researchers, modeling and predicting shocks tests the limits of both inductive bias and network expressivity, especially in generalization across flow regimes. By including shock-resolving simulations in our dataset, we offer a benchmark that pushes beyond existing datasets constrained to subsonic, smooth flows—highlighting scenarios where naive ML architectures fail and physics-aware models are essential.
@@ 1.3 Main contribution
To summarize and contrast the key new additions the current dataset brings to the literature, we refer the readers to Table˜1 and a summary of the input fields can be found in Table˜2.
@@ 2.1 Geometry
Airfoil geometries can be efficiently represented in a reduced state space using modal decomposition techniques. In particular, a low-dimensional basis can be extracted from a dataset of classic airfoils via proper orthogonal decomposition (POD), enabling compact representation and efficient sampling of shape variations. Notably, airfoils designed for fully turbulent operation differ significantly in geometry—and consequently in modal characteristics—from those optimized to exploit laminar–turbulent transition mechanisms. Thus, we create two airfoil shape datasets for the fully turbulent (FT) airfoils and natural laminar flow (NLF) airfoils. In the current dataset, we sampled 30,000 FT and 4,800 NLF airfoils. An example of this sampling is shown in Figure˜2 for the FT airfoils. We refer the readers to Li et al. (2019) for further information on sampling large datasets of geometric variations in a particular class of airfoils.
@@ 2.2 Mesh generation
Mesh generation for the airfoil simulations was carried out using pyHyp by Secco et al. (2021) [^4] https://github.com/mdolab/pyhyp. pyHyp generates high-quality O-type multi-block meshes around airfoils using a hyperbolic marching method, which extrudes mesh lines from an initial surface mesh by solving a system of hyperbolic partial differential equations (PDEs) that control cell spacing, orthogonality, and smoothness. One mesh generated by pyHyp is shown in Figure˜3. This approach enables accurate boundary layer resolution and consistent mesh quality, making it well suited for RANS solvers such as ADflow. The method is fully automated and robust across large geometric datasets, supporting high-throughput CFD simulations.
@@ 2.3 Solver and numerical considerations
All simulations were steady state and performed on a far-field mesh with the airfoil body-fitted at the center of the mesh, as shown in Figure˜3. The wall was given a no-slip boundary condition and the outermost edge of the computational domain was given a far-field boundary condition. Further details on the numerical methods employed in data generation are elaborated in the supplementary data.
An overview of the flow field figures and flow regimes in the dataset are summarized in Figure˜1. In the figure, $C_{p}$ is the coefficient of pressure and $\mu_{t}/\mu$ is the ratio of the turbulent eddy viscosity to the fluid’s molecular viscosity.
@@ 3 Dataset description
We sample a wide range of Mach numbers, Reynolds numbers, and AoAs. We have the largest ranges of Mach numbers and Reynolds numbers among all datasets Table˜1. The range of AoA is narrower than in many other datasets, as large AoAs are typically avoided in practice to maintain aerodynamic efficiency. In addition, with a larger AoA, RANS model is no longer able to capture the flow field accurately due to flow separation. For each airfoil geometry, we conduct around $13$ simulations with different Mach, Re, and AoA. These three input variables were sampled using a Lattice HyperCube Sampling approach.
The dataset is composed of three parts: The majority of the simulations are done using the fully turbulent flow assumption with SA model together with fully turbulent airfoil geometry class (more details see Section˜2). This dataset represents the most commonly used operation condition. The second dataset also uses the fully turbulent assumption but uses the NLF geometry class. Finally, the last dataset has a one-to-one correspondance with the second dataset with identical flight conditions and geometries. The only difference is that the last dataset employs an $e^{N}$-based transition model to capture the laminar–turbulent transition, whereas the second dataset assumes fully turbulent flow from the leading edge of the airfoil.
@@
Turbulence Model Airfoil Class No. of Airfoils Mach Re ($\times 10^{6}$) AoA () Net Simulations
SA model FT 30,000 $[0.1,0.85]$ $[1,10]$ [2,6][-2,6] 400,000
NLF 4,800 $[0.1,0.85]$ $[1,10]$ [2,6][-2,6] 50,000
$e^{N}$ transition model (+ trip term) NLF 4,800 $[0.1,0.85]$ $[1,10]$ [2,6][-2,6] 50,000
@@ 3.1 Dataset access and contents summary
The dataset is publicly available under the CC-BY-SA license at the Harvard Dataverse Kanchi et al. (2025). The dataset contents are summarized in Table˜3. All data processing and visualization scripts will be released in a public repository referenced in the supplementary material. The CGNS files listed in Table˜3 have the coefficient of pressure, velocity vector data and the mach number. Each CGNS file retains the grid and the values of the computed flow variables in that grid. For the NLF-transition simulations, the CGNS files have additional information, such as turbulence intermittency, temperature, skin-friction coefficients and density. The lift and drag coefficients ($C_{L}$ and $C_{D}$) are placed in the tabular file as described in Table˜3.
@@
airfoil_geometry/ Airfoil coordinates and generation code (in supplementary repo2).
@@ 3.3 Limitations of the Dataset
4 Preliminary results for ML
In this section, we present the key ML results using UniFoil. We describe the training architecture and present the predicted results bench–marked against results from ADflow. More details on data pre–processing and training and testing metrics can be found in Appendix˜D.
4.1 Neural network architecture
The neural network trained on the UniFoil dataset is an encoder-decoder model with a latent space deep neural network (DNN). An overview of the architecture of the neural network is shown in Figure˜4. A summary of the training setup is shown in Table˜4. In the table the “Param #” field represents the quantity of variables that can be modified during the training procedure. Densely connected layers such as the ones prevalent in the DNN have a large number of these parameters as all 6134 nodes are connected to every single node below. This model is trained on the pressure-coefficient data extracted from each airfoil simulation. The model is split into two main components with different layer designs used to down-sample and up-sample the input space. Further details on the exact architecture specifics and data pre–processing step are presented in Sections˜D.1 and D.2.
Figure 4: An overview of the entire training pipeline. The modal coefficients $a_{i}$ are obtained from Equation˜5.